A mass of 2kg of water is contained in a vessel having a volume of 0.2m^3 at 4℃. Determine the density and the specific volume of water at the given temp.

1
Expert's answer
2022-05-23T16:24:37-0400

"\rho=\frac mV=10~\frac{kg}{m^3},"

"v=\frac 1\rho=0.1~\frac{m^3}{kg}."